#5 EllEN WISHES TO MIX CANDY WORTH $1.85 PER POUND WITH CANDY WORTH $3.33 PER POUND TO FORM 29 POUNDS OF A MIXTURE WORTH $2.62 PER POUND. HOW MANY POUNDS OF THE MORE EXPENSIVE CANDY SHOULD SHE USE?
Answer by stanbon(48510)

EllEN WISHES TO MIX CANDY WORTH $1.85 PER POUND WITH CANDY WORTH $3.33 PER POUND TO FORM 29 POUNDS OF A MIXTURE WORTH $2.62 PER POUND. HOW MANY POUNDS OF THE MORE EXPENSIVE CANDY SHOULD SHE USE?
----------------------------------
Equation:
value + value = value
1.85x + 3.33(29-x) = 2.62*29
Multiply thru by 100 to get:
185x + 333*29 - 333x = 262*29
-148x = -71*29
---
x = 13.91 lb (amount of $1.85 candy needed in the mix)
--------------
29-x = 29-13.91 = 15.09 lb (amount of $3.33 candy needed in the mix)
========================
Cheers,
Stan H.